OpenStudy (anonymous):

What are the zeros of the polynomial function: f(x) = x3 – 5x2 + 6x ?

OpenStudy (anonymous):

x(x^2-5x+6)=x(x-6)(x+1) x=0,6,-1

OpenStudy (lalaly):

\[x(x-3)(x-2)\]

OpenStudy (lalaly):

x=0, x=3 , x=2

OpenStudy (anonymous):

ohh

OpenStudy (anonymous):

:|

OpenStudy (anonymous):

ah thank you lalaly

OpenStudy (anonymous):

was going a little too fast

OpenStudy (lalaly):

=)

OpenStudy (anonymous):

beautiful helpful people